# Synapse: Learning Preferential Concepts from Visual Demonstrations

### v1

@@ 1 Introduction
Figure 1: Illustration of Synapse for identifying good contingency locations. (a) Synapse learns a neuro-symbolic program that represents the preferential concept based on user demonstrations, which include both a NL explanation and a physical demonstration. The learning algorithm consists of three steps, namely updating the concept library, synthesizing a program sketch, and performing parameter synthesis; (b) The new sketch gets synthesized based on the previous sketch and the current natural language description; (c) Parameter synthesis determines suitable values for numeric parameters in the program sketch; (d) Synapse evaluates the program on a new query image to return a preference (in this case, boolean) mask over the input image.
@@
Imagine trying to come up with a definition of “a good taxi drop-off location”. One person may consider a spot to be a good drop-off location depending on whether it is close to the door of a building, while someone else might want it in the shade. Such concepts vary from person to person and inherently depend on their preferences. We call them preferential concepts, and we are interested in the problem of preference learning from visual input. Learning preferences is important because we want systems that are customizable and that can adapt to end-users. This problem is quite related to the task of visual concept learning, wherein much of the work focuses on learning concepts such as having the color red or being to the left of another object [5, 52, 40, 43, 35, 23, 33, 61, 13, 41, 44, 31, 49, 26, 55, 25]. All such prior work assumes there is a ground-truth for the concept, i.e., the definition of the concept does not differ among people, and as a consequence, sufficiently many examples are available, and can be objectively evaluated. We refer to such concepts as factual concepts. While most prior work that learns visual concepts exploits the availability of large datasets such as CLEVR [29], those methods cannot be applied to preference learning because it is a data-impoverished setting by its very nature: a single individual can put up with providing only so much data. This limitation is also present in most of the preference learning work in the reinforcement learning literature as well, where human preferences are represented as neural networks [59, 10] or latent reward models [60, 2, 14, 48, 7, 15]. Furthermore, because preferences are inherently individual, they can depend on entirely different concepts, such as in the drop-off location example above (i.e., based on proximity to door as opposed to being in the shade). This requires learning novel visual concepts in a hierarchical manner. Lastly, coming up with a complete definition of a preferential concept at once is itself a hard problem: it is much easier for someone to show examples that satisfy their intuition as humans tend to build their notion of a preferential concept over time. Thus, preference learning calls for an approach that can handle incremental learning from visual demonstrations.
@@ 3.1 Representing Preferences
{mathpar} InputsqQoO
{mathpar} Constantsv{Int, Real, …}pP
{mathpar} Termst := q  ∣ o  ∣ v∣ f(t1, …, tn)   where f ∈$\mathcal{C}$f
[…]
{mathpar} Programsπ:=p  ∣ if  (ϕthen  π else  π
@@ 4.1 Implementation Details
Models. We use Grounded-SAM [46, 39, 32] zero-shot VLM for object detection and a custom terrain model with SegFormer architecture [62] trained to segment terrains from sensor data. We use both of these models to get segmentation masks of the neural concepts (i.e., objects and terrains) currently in the concept library. We use GPT-4 [1] as the language model for sketch synthesis in Synapse.
@@ 4.2 Results
To evaluate Synapse, we create a dataset of 815 labeled images taken from the UT Austin campus area, where the labels mark the locations on the images that are consistent with the intended user preference for each of the three tasks. We split the dataset into three sets: train, in-distribution test, and out-of-distribution test sets. The train and in-distribution sets belong to the same part of the UT Austin campus region, while the out-of-distribution set belongs to a different part of the region. Table 1 shows the comparison against various baselines. We use mean Intersection-Over-Union (mIOU) as the metric and evaluate the following baselines: (1) pure neural models based on SegFormer [62] architecture with pretrained weights, with and without depth input, fine-tuned on our custom dataset; (2) GPT4 [1] with vision capabilities, and (3) VisProg [22]. The ‘+’ for the latter two indicate additional prompting that provides additional information about the underlying reasoning behind the preferential concept. More details about the dataset and how different baselines are trained and/or evaluated can be found in the supplementary material.
@@ 4.3 Ablations
We investigate three types of ablations: (1) NN-ablations, in which we compare the performance of neural baselines against Synapse when trained on the same number of samples (i.e., 29), (2) LLM-based ablations, where we see how using different LLMs affect the performance of Synapse, and (3) framework ablations where we test different design choices. We investigate ablations based on these framework features: (1) feat1: whether it queries the user for auxiliary concepts, (2) feat2: whether it performs lifelong learning by building on its concept library, and (3) feat3: whether it has got direct access to the full concept library learned by Synapse. More details are provided in the supplementary material.
@@
feat1 feat2 feat3 LLM mIOU (%)
Synapse GPT-4 76.11
Synapse-SynthDirect GPT-4 60.74
Synapse-SynthDirect+ GPT-4 68.86
Synapse-SynthCaP GPT-4 64.11
Synapse-CodeLLama CodeLLama [47] 69.88
Synapse-StarCoder StarCoder [34] 63.62
Synapse-PaLM2 PaLM2 [6] 71.62
SF-RGB-b0 - - - - 54.58
SF-RGB-b5 - - - - 56.30
SF-RGBD-b0 - - - - 55.84
SF-RGBD-b5 - - - - 63.81
@@ 0.B.1 Baselines
Tables 5, 6 and 7 show the full evaluation for the three concepts, with the IOU for each class. Due to unavoidable data imbalance on the pixel-level (i.e., $\approx$ 90% negative pixels against 10% positive class pixels), it is much easier to learn the negative class.
@@ 0.B.2 Ablations
    LLM-ablations: We test the performance of our overall framework using other language models available: (1) CodeLLama [47], (2) StarCoder [34], and (3) PaLM2 [6]. We see that it is essential for the LLM to be strong enough for the performance of Synapse to be good.
The results reaffirm the design choices made in Synapse.

### v2

@@ 1 Introduction
Figure 1: Illustration of Synapse for identifying good taxi drop-off locations. (a) Synapse learns a neuro-symbolic program that represents the preferential concept based on user demonstrations, which include both a NL explanation and a physical demonstration. The learning algorithm consists of three steps, namely updating the concept library, synthesizing a program sketch, and performing parameter synthesis; (b) Synapse evaluates the program on a new query image to return a preference (in this case, boolean) mask over the input image; (c) given a user demonstration, (d) we extract entities of relevance from the NL input; (e) for the new observed predicates, Synapse recursively calls itself to first learn these auxiliary predicates from new user demonstrations (here, in the example, it takes the form of NL-only demonstrations) followed by (f) code generation; (g) once all the newly seen predicates are learned, a compact form representation (i.e., CNF) is extracted from the NL input, and (h) partial evaluations are carried out from the physical demonstrations; (i) using the previous program sketch and the current CNF, a new sketch is synthesized; (j) finally, parameter synthesis fills the holes in the sketch using the partial evaluations.
@@
Imagine trying to come up with a definition of “a good taxi drop-off location”. One person may consider a spot to be a good drop-off location depending on whether it is close to the door of a building, while someone else might want it in the shade. Such concepts vary from person to person and inherently depend on their preferences. We call them preferential concepts, and we are interested in the problem of preference learning from visual input. Learning preferences is important because we want systems that are customizable and that can adapt to end-users (e.g., consider a scenario where a user buys a home robot). This problem is quite related to the task of visual concept learning, wherein much of the work focuses on learning concepts such as having the color red or being to the left of another object [5, 53, 40, 44, 35, 23, 33, 62, 13, 41, 45, 31, 50, 26, 56, 25]. All such prior work assumes there is a ground-truth for the concept, i.e., the definition of the concept does not differ among people, and as a consequence, sufficiently many examples are available, and can be objectively evaluated. We refer to such concepts as factual concepts. While most prior work that learns visual concepts exploits the availability of large datasets such as CLEVR [29], those methods cannot be applied to preference learning because it is a data-impoverished setting by its very nature: a single individual can put up with providing only so much data. This limitation is also present in most of the preference learning work in the reinforcement learning literature as well, where human preferences are represented as neural networks [60, 10] or latent reward models [61, 2, 14, 49, 7, 15]. Furthermore, because preferences are inherently individual, they can depend on entirely different concepts, such as in the drop-off location example above (i.e., based on proximity to door as opposed to being in the shade). This requires learning novel visual concepts in a hierarchical manner. Lastly, coming up with a complete definition of a preferential concept at once is itself a hard problem: it is much easier for someone to show examples that satisfy their intuition as humans tend to build their notion of a preferential concept over time. Thus, preference learning calls for an approach that can handle incremental learning from visual demonstrations.
@@ 3.1 Representing Preferences
{mathpar}
Inputs
qQoO
{mathpar}
Constants
v{Int, Real, …}pP
{mathpar}
Terms
t := q  ∣ o  ∣ v∣ f(t_1, …, t_n)   where f ∈$\mathcal{C}$_f
{mathpar}
Conditions
ϕ:=p_c(t_1, …, t_n) where p_c ∈$\mathcal{C}$_b∣ ¬ϕ ∣ ϕϕ ∣ ϕϕ
{mathpar}
Programs
π:=p  ∣ if  (ϕthen  π else  π
@@ 4.1 Implementation Details
Models. We use Grounded-SAM [47, 39, 32] zero-shot VLM for object detection and a custom terrain model with SegFormer architecture [63] trained to segment terrains from sensor data, since we observed that open-set visual models perform pretty poorly on terrains. We use both of these models to get segmentation masks of the neural concepts (i.e., objects and terrains) currently in the concept library. We use GPT-4 [1] as the language model for sketch synthesis in Synapse.
@@ 4.2 Results
To evaluate Synapse, we create a dataset of 815 labeled images taken from the UT Austin campus area, where the labels mark the locations on the images that are consistent with the intended user preference for each of the three tasks. We split the dataset into three sets: train, in-distribution test, and out-of-distribution test sets. The train and in-distribution sets belong to the same part of the UT Austin campus region, while the out-of-distribution set belongs to a different part of the region. Table 1 shows the comparison against various baselines. We use mean Intersection-Over-Union (mIOU) as the metric and evaluate the following baselines: (1) pure neural models based on SegFormer [63] architecture with pretrained weights, with and without depth input, fine-tuned on our custom dataset; (2) GPT4 [1] with vision capabilities, and (3) VisProg [22]. The ‘+’ for the latter two indicate additional prompting that provides additional information about the underlying reasoning behind the preferential concept. More details about the dataset, how different baselines are trained and/or evaluated, and results with error bounds can be found in the supplementary material.
@@ 4.3 Ablations
We investigate four types of ablations: (1) NN-ablations, in which we compare the performance of neural baselines against Synapse when trained on the same number of samples (i.e., 29), (2) LLM-based ablations, where we see how using different LLMs affect the performance of Synapse, (3) VLM-based methods where we test different vision-language models, and (4) framework ablations where we test different design choices. We investigate ablations based on these framework features: (1) feat1: whether it queries the user for auxiliary concepts, (2) feat2: whether it performs lifelong learning by building on its concept library, and (3) feat3: whether it has got direct access to the full concept library learned by Synapse. More details are provided in the supplementary material.
@@
feat1 feat2 feat3 LLM VLM mIOU (%)
Synapse GPT-4 [1] DINO-SAM [47] 76.11
Synapse-SynthDirect GPT-4 [1] DINO-SAM [47] 60.74
Synapse-SynthDirect+ GPT-4 [1] DINO-SAM [47] 68.86
Synapse-SynthCaP GPT-4 [1] DINO-SAM [47] 64.11
Synapse-CodeLLama CodeLLama [48] DINO-SAM [47] 69.88
Synapse-StarCoder StarCoder [34] DINO-SAM [47] 63.62
Synapse-PaLM2 PaLM2 [6] DINO-SAM [47] 71.62
Synapse-OWLViTSAM GPT-4 [1] OWLViT [43] 70.17
Synapse-GroupViT GPT-4 [1] GroupViT [64] 73.41
SF-RGB-b0 - - - - - 54.58
SF-RGB-b5 - - - - - 56.30
SF-RGBD-b0 - - - - - 55.84
SF-RGBD-b5 - - - - - 63.81
@@ Appendix 0.A Implementation Details
Figure 7: Results for the CONTINGENCY concept
Figure 9: Representative samples showing differences in multi-user preferences
@@ 0.B.1 Baselines
Tables 5, 6 and 7 show the full evaluation for the three concepts, with the IOU for each class. Due to unavoidable data imbalance on the pixel-level (i.e., $\approx$ 90% negative pixels against 10% positive class pixels), it is much easier to learn the negative class. Figure 7(a) shows the results with error bounds (across five runs) for the CONTINGENCY concept.
@@ 0.B.2 Ablations
    LLM and VLM ablations: We test the performance of our overall framework using some of the other language models: (1) CodeLLama [48], (2) StarCoder [34], and (3) PaLM2 [6], and a few vision-language models: (1) OWL-ViT [43] with SAM [32], and (2) GroupViT [64]. We see that the performance of Synapse depends a fair bit on the strength of the underlying foundation models.
Figure 7(b) shows the results (with error bounds, across five runs) that reaffirm the design choices made in Synapse.